EDITORIAL

Terry Goodman

In recent years, many mathematics teachers have had their students celebrate March 14 (3/14) as Pi Day. In Missouri, the Missouri Council of Teachers of Mathematics has taken the lead in encouraging mathematics teachers to incorporate Pi Day activities into their classrooms. MCTM has even provided activities and materials for teachers to use with their students (further information can be found at: \underline{http://www.MoCTM.org/PiDay/PiDayActivities.html}). Feedback from teachers and students has indicated that Pi Day has been a great success and that both groups have found the activities of the day to be interesting and informative.

One of our readers, Jeganathan Sriskandarajah has provided MJMS with information concerning the most recent celebration of Pi Day in Wisconsin. Below is a copy of a letter that was sent by the math club for which he serves as advisor.

Faculty Advisor/Math Club Pi Day Activities

The Math Club invites you to participate in the celebration of Pi Day, scheduled for Friday, March 14. Coincidentally, the date is also the birth date of Albert Einstein. As most people know, Pi, the mathematical constant, is the ratio of a circle’s circumference to its diameter, and numerically equal to approximately 3.14. It has inspired poems, limericks, cartoons, and numerous jokes. Both Archimedes and da Vinci seemed to have been obsessed with it. It has been calculated to eight billion digits using super computers.

We will celebrate this day with the following activities:

1. A math competition for the two–year colleges of Wisconsin....

2. A presentation, “The Calculation of Pi,” by Professor Ranjan Roy of Beloit College.

3. A pie–eating contest, the “Pie-athlon”: the winner will be determined by the number of pies eaten in 3 minutes and 14 seconds! There will be prizes for both men and women champions.

4. The Pi Day Poster Competition encourages students to discover the usefulness and significance of Pi in creative ways. Prizes will be awarded to winners and their instructors. In submitting a poster entry, students agree that the poster may be displayed at the MATC math club presentations, included on its web site and featured in the publication “Pi in the sky”....

You can find further information on this club’s celebration of Pi Day posted on: \underline{http://matcmadison.edu/is/as/math/mathclub/Piday03/Piday03.html}.